OpenStudy (anonymous):

Find the function of the series sum (x^2 +1)^2k

OpenStudy (anonymous):

we do this in steps, and of course the answer will only be valid for \(|x^2+1|^2<1\)

OpenStudy (anonymous):

trick is to make it look like the geometric series \[\sum x^k=\frac{1}{1-x}\] or rather to use that as a starting point

OpenStudy (anonymous):

what is the geon. series...

OpenStudy (anonymous):

then \[\sum x^{2k}=\frac{1}{1-x^2}\] by replacing \(x\) by \(x^2\)

OpenStudy (anonymous):

hold the phone is this line ok? \[\sum x^k=\frac{1}{1-x}\]

OpenStudy (anonymous):

because that is our starting point. if that is not clear, then we cannot continue if that is clear, then it is algebra from here on in

OpenStudy (anonymous):

but keep i mind at this is \[1\div1+x ^{2}\] not the 1/1-x...

OpenStudy (anonymous):

we have this \(\sum (x^2+1)^2\) right?

OpenStudy (anonymous):

\[\sum_{}^{} (x ^{2}+1)^{2k}\]

OpenStudy (anonymous):

so we are going to make some replacements in this identity \[\sum x^k=\frac{1}{1-x}\] first we find \[\sum x^{2k}=\frac{1}{1-x^2}\] by replacing \(x\) by \(x^2\) then we replace \(x\) in the above by \((x^2+1)\) to get \[\sum (x^2+1)^{2k}=\frac{1}{1-(x^2+1)^2}\]

OpenStudy (anonymous):

then maybe a little alegra to clean up the denominator and write \[\sum (x^2+1)^{2k}=\frac{1}{x^2(x^2+2)}\]

OpenStudy (anonymous):

*algebra